OpenStudy (anonymous):

fill in the blank. y^2 + 18x + _____=

OpenStudy (anonymous):

= what? o.o

OpenStudy (anonymous):

sorry, that's where I factor once i fill in the blank. ignore the = sign

OpenStudy (anonymous):

ohs /o.o\

OpenStudy (anonymous):

the directions are: fill in the blank so that each expression is a trinomial square; factor

OpenStudy (anonymous):

she doesn't know maybe

OpenStudy (dtan5457):

To get the value of C (your missing value) on a perfect square trinomial, divide the value of B by 2, and square that value. So y^2+18x+c 18=b 9^2=81 c=81 (x+9)(x+9)<<<perfect square factor trinomial

OpenStudy (anonymous):

Great! thanks!

OpenStudy (dtan5457):

np feel free to medal and fan and tag me into more questions

OpenStudy (anonymous):

This is my first time using this, I don't know how to do what you said.